# The DESI DR1 Peculiar Velocity Survey: Growth Rate Measurements from the Galaxy Power Spectrum

### v1

@@
The large-scale structure of the Universe and its evolution encapsulate a wealth of cosmological information. A powerful means of unlocking this knowledge lies in measuring the auto-power spectrum and/or the cross-power spectrum of the galaxy density and momentum fields, followed by the estimation of cosmological parameters based on these spectrum measurements. In this study, we generalize the cross-power spectrum model to accommodate scenarios where the density and momentum fields are derived from distinct galaxy surveys. The growth rate of the large-scale structures of the Universe, commonly represented as $f\sigma_{8}$, is extracted by jointly fitting the monopole and quadrupole moments of the auto-density power spectrum, the monopole of the auto-momentum power spectrum, and the dipole of the cross-power spectrum. Our estimators, theoretical models and parameter-fitting framework have been tested using mocks, confirming their robustness and accuracy in retrieving the fiducial growth rate from simulation. These techniques are then applied to analyze the power spectrum of the DESI Bright Galaxy Survey and Peculiar Velocity Survey, and the fit result of the growth rate is $f\sigma_{8}=0.440^{+0.080}_{-0.096}$ at effective redshift $z_{\rm eff}=0.07$. By synthesizing the fitting outcomes from correlation functions, maximum likelihood estimation and power spectrum, yields a consensus value of $f\sigma_{8}(z_{\rm eff}=0.07)=0.450^{+0.055}_{-0.055}$, and correspondingly we obtain $\gamma=0.580^{+0.110}_{-0.110}$, $\Omega_{\mathrm{m}}=0.301^{+0.011}_{-0.011}$ and $\sigma_{8}=0.834^{+0.032}_{-0.032}$. The measured $f\sigma_{8}$ and $\gamma$ are consistent with the prediction of the $\Lambda$ Cold Dark Matter Model and General Relativity.
@@ 1 Introduction
An approach to assess the cosmological models involves measuring the low-order kinematic moments of the cosmic flow field based on the radial components of galaxies’ peculiar velocities, and subsequently comparing these estimates with the predictions of the cosmological models (Kaiser, 1988; Strauss & Willick, 1995; Feldman et al., 2010; Scrimgeour et al., 2016; Zhang et al., 2017; Qin et al., 2018, 2019b; Qin, 2021; Qin et al., 2021; Whitford et al., 2023). However, due to the significant cosmic variance inherent in low-redshift velocity surveys, the constraints imposed on cosmological models through this method are typically not tight.
@@
One technique involves the reconstruction of the cosmological density and velocity fields (Nusser et al., 1991; Kitaura et al., 2012; Wang et al., 2012), or the cosmography of the local Universe (Courtois et al., 2013; Dupuy & Courtois, 2023). The parameters $f\sigma_{8}$ and $b\sigma_{8}$ can be estimated by comparing the observed peculiar velocities of galaxies with the reconstructed velocity field (Carrick et al., 2015; Said et al., 2020; Boruah et al., 2022). However, systematic errors associated with this method are difficult to quantify, and previous studies have indicated that the errors tend to be underestimated (Turner & Blake, 2023; Blake & Turner, 2024). Recently, artificial intelligence techniques have been increasingly applied to field reconstruction (Hong et al., 2021; Qin et al., 2023a; Wu et al., 2023), offering novel opportunities in this research area. Moreover, since the covariance matrix of the density and velocity fields can be formulated in terms of $f\sigma_{8}$ and $b\sigma_{8}$, these parameters can alternatively be estimated by maximizing the likelihood function under the assumption of a Gaussian distribution (Johnson et al., 2014; Adams & Blake, 2017; Lai et al., 2023; Carreres et al., 2023; Rocher et al., 2023; Ravoux et al., 2025). However, this method remains computationally demanding. Another commonly used approach is the two-point statistic method, which has played a significant role in cosmology over the past decades. Based on this method, $f\sigma_{8}$ and $b\sigma_{8}$ can be derived from the galaxy two-point correlation functions, the velocity-correlation functions and the galaxy-velocity cross-correlation functions (Gorski et al., 1989; Howlett et al., 2015; Dupuy et al., 2019; Wang et al., 2018; Turner et al., 2021, 2023; Qin et al., 2022, 2023b; Lyall et al., 2023). Nevertheless, these approaches tend to be computationally intensive too.
Instead of studying the two-point correlations in configuration space, we can also analyze them in Fourier space, i.e. analyzing the so-called power spectrum. The concept of the cosmological power spectrum was first introduced by Kaiser (1987). The first measurement of the (galaxy) density power spectrum was carried out by Kaiser & Peacock (1991). The methodology for estimating the density power spectrum was later refined by Feldman et al. (1994), forming the standard approach widely used today. In addition to these foundational contributions, Yamamoto et al. (2006) and Bianchi et al. (2015) further developed estimators for the multipoles of the density power spectrum.
@@
We constrain $f\sigma_{8}$ using Data Release 1 (DR1) from the Bright Galaxy Survey (BGS, Hahn et al. 2023) and Peculiar Velocity Survey (DESI-PV, Carr et al. in prep.; Ross et al. in prep.; Douglass et al. in prep.) of the Dark Energy Spectroscopic Instrument (DESI, DESI Collaboration et al. 2025a, b). The DESI-PV DR1 stands as the largest peculiar velocity survey ever assembled to date. This paper forms part of a comprehensive series results from DESI-PV DR1 (Turner et al., in prep.; Lai et al., in prep.; Ross et al., in prep.; Douglass et al., in prep.; Bautista et al., in prep.; Carr et al., in prep.; Nguyen et al., in prep.).
@@ 2.1.1 The BGS data
The Bright Galaxy Survey (BGS, Hahn et al. 2023) Bright sample has undergone refinement by Turner et al. (in prep.), who introduced tailored selection criteria to enhance its suitability for analyses involving peculiar velocities. The final BGS data catalogue employed in our study encompass a total of 415,523 galaxies, with an r-band absolute magnitude threshold set at -17.7. The sky distribution of these galaxies is illustrated by the black dots in Fig.1. Furthermore, the redshifts of these galaxies, measured in the CMB frame, span the range from $z=0.01$ to $z=0.1$, as depicted by the black bars in Fig.2. The galaxy mean number density $\bar{n}({\bf r})$ is shown in the top-left panel of Fig.3. The calculation of $\bar{n}({\bf r})$ is detailed in Bautista et al. (in prep.). In this paper, the density field is derived from the BGS catalogue.
@@ 2.1.2 The DESI-PV data
Due to the peculiar motions of galaxies, the apparent distance of a galaxy, denoted as $d_{z}$, deviates from its true comoving distance, $d_{h}$. This discrepancy is measureable and can be quantified through the logarithmic distance ratio, defined as $\eta\equiv\log_{10}\frac{d_{z}}{d_{h}}$. It can be converted to the line-of-sight peculiar velocity using the velocity estimator Eq.57. Within the DESI-PV survey, two distinct populations of galaxies are observed: spiral galaxies, also known as late-type galaxies, and elliptical galaxies, referred to as early-type galaxies. The $\eta$ for late-type galaxies is derived from the Tully-Fisher (TF, Tully & Fisher 1977) relation, whereas for early-type galaxies, it is inferred from the Fundamental Plane (FP, Djorgovski & Davis 1987, Dressler et al. 1987) relation. In DESI-PV, the Tully-Fisher sample (Douglass et al., in prep.) comprises a total of 6,806 late-type galaxies. The sky distribution of these galaxies is illustrated by the blue dots in Fig.1. Furthermore, the redshifts of these galaxies span the range from $z=0.015$ to $z=0.1$, as depicted by the blue bars in Fig.2. In contrast, the Fundamental Plane sample (Ross et al., in prep.) consists of 73,822 early-type galaxies. Their sky distribution is marked by gray dots in Fig.1, and their redshifts span the interval from $z=0.01$ to $z=0.1$, as shown in the gray bars of Fig.2.
@@ 2.2 Mocks
A considerable number of mocks is indispensable for validating the algorithm and accurately estimating the uncertainties in the power spectrum and associated parameters. Bautista et al. (in prep.) offers a comprehensive set of 675 BGS mocks, meticulously designed to mirror the characteristics of the BGS data. In parallel, 675 FP mocks have been constructed to emulate the FP data, alongside 675 TF mocks tailored to align with the TF data. Each set of BGS, FP, and TF mocks shares the same observer, resulting in a total of 675 mock sets. These mocks are derived from the AbacusSummit simulation, and the fiducial value of $f\sigma_{8}$ for these mocks is 0.466 at an effective redshift of $z=0.2$.
@@ 2.3 The random catalogs
The random catalogues for both the data and mocks are also supplied by Bautista et al. (in prep.). In order to precisely extract cosmological parameters from the power spectrum, it is imperative to appropriately account for the survey geometry of both BGS and DESI-PV. Employing fast Fourier transforms to measure the power spectrum in the context of a non-periodic and incomplete survey geometry leads to a convolution of the true power spectrum with the survey’s window function. To address this, The random catalogue will be employed to estimate the window function, thereby enabling the application of an equivalent convolution to our theoretical model during the data fitting process. The random catalogues for FP and TF are generated independently, with the objective of more faithfully replicating the distinct observational features of the TF and FP surveys, respectively. The random catalogue for DESI-PV is constructed by combining the random catalogues of TF and FP.
@@ 3.1 The Field Functions
We favor the momentum power spectrum over the velocity power spectrum because, in principle, the velocity field is a continuous entity defined everywhere in space—even in regions devoid of galaxies. However, our measurements are limited to locations where galaxies (and thus mass) are present, effectively yielding a mass-weighted, or momentum-based power spectrum. In regions devoid of galaxies, we record zero velocity not because the velocity is truly zero, but due to the absence of tracers. Additionally, the momentum power spectrum retains the same fundamental information as the velocity power spectrum, and with additional non-linear contributions introduced by the $(1+\delta)$ term that be incorporated into the theoretical modeling.
@@
which ensures that the amplitude of the measured power aligns with it in a universe unaffected by survey selection effects. The computation of the mean galaxy number density $\bar{n}_{\delta}({\bf r})$ at position $\bf r$} is detailed in Bautista et al. (in prep.) or see Fig.3. As indicated in Eq.6, to estimate the density field, it is necessary to subtract the random catalog, this is because we measure the density contrast field $\delta$, which can only be accomplished if we possess knowledge of, and are able to subtract, the mean number of galaxies within each grid cell which estimated from randoms of that cell.
@@
which again adjusts the amplitude of the measured power to conform with it in an idealized universe without survey selection biases. The computation of the mean galaxy number density $\bar{n}_{p}({\bf r})$ at position $\bf r$} is also detailed in Bautista et al. (in prep.) or see Fig.3. Notably, no random catalog is subtracted when estimating the momentum field. This is because that the mean velocity at any spatial point is inherently zero; consequently, subtracting a random catalogue during the construction of the momentum field would be equivalent to subtracting zero.
@@ 3.2 The estimators of the power spectrum multipoles
As outlined in Appendix G.1 (or see Paper III), the estimator for the cross-power spectrum multipoles is defined through  [^3] There is a small typos in Equation 19 of Paper III, while the code used in Paper III is correct.
@@
where $\ell=0,1,2,3,4,$.... Consequently, based on Eq.13 or Eq.14, the corresponding even-multipoles of the auto-power spectrum is expressed in (see Appendix G.3 for more discussion)
@@
where $d\Omega_{k}$ denotes the solid angle. See Appendix D for the expressions of $T_{0}$, $T_{2}$ and $T_{4}$. It is important to note that, unlike in Bianchi et al. (2015), the normalization factors $A_{\delta}$ and $A_{p}$ are inherently incorporated within the field functions in our formulation, and our definition of Fourier transformation diverges from them too. For multipoles with $\ell>0$, the shot noise becomes negligible and can be safely ignored (Blake et al., 2010; Howlett, 2019). If ignoring the wide-angle or general relativistic effects, only the even-multipoles of the auto-power spectrum are non-zero (Bianchi et al., 2015; Beutler & Di Dio, 2020).

### v2

@@
The large-scale structure of the Universe and its evolution encapsulate a wealth of cosmological information. A powerful means of unlocking this knowledge lies in measuring the auto-power spectrum and/or the cross-power spectrum of the galaxy density and momentum fields, followed by the estimation of cosmological parameters based on these spectrum measurements. In this study, we generalize the cross-power spectrum model to accommodate scenarios where the density and momentum fields are derived from distinct galaxy surveys. The growth rate of the large-scale structures of the Universe, commonly represented as $f\sigma_{8}$, is extracted by jointly fitting the monopole and quadrupole moments of the auto-density power spectrum, the monopole of the auto-momentum power spectrum, and the dipole of the cross-power spectrum. Our estimators, theoretical models and parameter-fitting framework have been tested using mocks, confirming their robustness and accuracy in retrieving the fiducial growth rate from simulation. These techniques are then applied to analyze the power spectrum of the DESI Bright Galaxy Survey and Peculiar Velocity Survey, and the fit result of the growth rate is $f\sigma_{8}=0.440^{+0.080}_{-0.096}$ at effective redshift $z_{\rm eff}=0.07$. By synthesizing the fitting outcomes from correlation functions, maximum likelihood estimation and power spectrum, yields a consensus value of $f\sigma_{8}(z_{\rm eff}=0.07)=0.450^{+0.055}_{-0.055}$, and correspondingly we obtain $\gamma=0.580^{+0.110}_{-0.110}$, $\Omega_{\mathrm{m}}=0.301^{+0.011}_{-0.011}$ and $\sigma_{8}=0.834^{+0.032}_{-0.032}$. The measured $f\sigma_{8}$ and $\gamma$ are consistent with the prediction of the $\Lambda$ Cold Dark Matter Model and General Relativity. The public code of power spectrum measurements, power spectrum theoretical models and window function convolution are available at https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum, and see Appendix I for more links. The examples for the utilization of the code can be found in this Jupyter Notebook https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/CosmPSPy/Code/Examp_PS.ipynb
@@ 1 Introduction
An approach to assess the cosmological models involves measuring the low-order kinematic moments of the cosmic flow field based on the radial components of galaxies’ peculiar velocities, and subsequently comparing these estimates with the predictions of the cosmological models (Kaiser, 1988; Jaffe & Kaiser, 1995; Strauss & Willick, 1995; Parnovsky et al., 2001; Watkins et al., 2009; Feldman et al., 2010; Hong et al., 2014; Nusser, 2014; Scrimgeour et al., 2016; Zhang et al., 2017; Peery et al., 2018; Qin et al., 2018, 2019b; Qin, 2021; Qin et al., 2021; Whitford et al., 2023; Watkins et al., 2023; Heinesen, 2023; Lopes et al., 2024). However, due to the significant cosmic variance inherent in low-redshift velocity surveys, the constraints imposed on cosmological models through this method are typically not tight.
@@
One technique involves the reconstruction of the cosmological density and velocity fields (Nusser et al., 1991; Zaroubi et al., 1995; Croft & Gaztanaga, 1997; Kudlicki et al., 2000; Branchini et al., 2002; Pike & Hudson, 2005; Erdoǧdu et al., 2006; Kitaura et al., 2012; Wang et al., 2012; Shi et al., 2018), or the cosmography of the local Universe (Courtois et al., 2013; Dupuy & Courtois, 2023). The parameters $f\sigma_{8}$ and $b\sigma_{8}$ can be estimated by comparing the observed peculiar velocities of galaxies with the reconstructed velocity field (Springob et al., 2014; Carrick et al., 2015; Ma et al., 2012; Said et al., 2020; Boruah et al., 2020, 2021, 2022; Hollinger & Hudson, 2024). However, systematic errors associated with this method are difficult to quantify, and previous studies have indicated that the errors tend to be underestimated (Turner & Blake, 2023; Blake & Turner, 2024a). Recently, artificial intelligence techniques have been increasingly applied to field reconstruction (Hong et al., 2021; Mao et al., 2021; Wu et al., 2021; Qin et al., 2023a; Ganeshaiah Veena et al., 2023; Wu et al., 2023; Lilow et al., 2024; Wang & Yang, 2024; Shi et al., 2025), offering novel opportunities in this research area. Moreover, since the covariance matrix of the density and velocity fields can be formulated in terms of $f\sigma_{8}$ and $b\sigma_{8}$, these parameters can alternatively be estimated by maximizing the likelihood function under the assumption of a Gaussian distribution (Johnson et al., 2014; Adams & Blake, 2017; Lai et al., 2023; Carreres et al., 2023; Rocher et al., 2023; Ravoux et al., 2025). However, this method remains computationally demanding. Another commonly used approach is the two-point statistic method, which has played a significant role in cosmology over the past decades. Based on this method, $f\sigma_{8}$ and $b\sigma_{8}$ can be derived from the galaxy two-point correlation functions, the velocity-correlation functions and the galaxy-velocity cross-correlation functions (Kaiser, 1984; Gorski et al., 1989; Howlett et al., 2015; Dupuy et al., 2019; Wang et al., 2018, 2021; Turner et al., 2021, 2023; Qin et al., 2022, 2023b; Lyall et al., 2023, 2024; Blake & Turner, 2024b; Nguyen et al., 2025; Turner et al., 2025). Nevertheless, these approaches tend to be computationally intensive too.
Instead of studying the two-point correlations in configuration space, we can also analyze them in Fourier space, i.e. analyzing the so-called power spectrum. The concept of the cosmological power spectrum was first introduced by Kaiser (1987). The first measurement of the (galaxy) density power spectrum was carried out by Kaiser & Peacock (1991). The methodology for estimating the density power spectrum was later refined by Feldman et al. (1994), forming the standard approach widely used today. In addition to these foundational contributions, Yamamoto et al. (2006) and Bianchi et al. (2015) further developed estimators for the multipoles of the density power spectrum. Other notable applications of the density power spectrum include studies by Yamamoto (2003); Sánchez & Cole (2008); Blake et al. (2010, 2011); Beutler et al. (2012); Johnson et al. (2014); Scoccimarro (2015); Blake et al. (2018); Beutler et al. (2019); Wang et al. (2019); Ivanov et al. (2020); Beutler & McDonald (2021).
@@
We constrain $f\sigma_{8}$ using Data Release 1 (DR1) from the Bright Galaxy Survey (BGS, Hahn et al. 2023) and Peculiar Velocity Survey (DESI-PV, Carr et al. 2025; Ross et al. 2025; Douglass et al. 2025) of the Dark Energy Spectroscopic Instrument (DESI, DESI Collaboration et al. 2025a, b). The DESI-PV DR1 stands as the largest peculiar velocity survey ever assembled to date. This paper forms part of a comprehensive series results from DESI-PV DR1 (Turner et al., 2025; Lai et al., 2025; Ross et al., 2025; Douglass et al., 2025; Bautista et al., 2025; Carr et al., 2025; Nguyen et al., 2025).
@@ 2.1.1 The BGS data
The Bright Galaxy Survey (BGS, Hahn et al. 2023) Bright sample has undergone refinement by Turner et al. (2025), who introduced tailored selection criteria to enhance its suitability for analyses involving peculiar velocities. The final BGS data catalogue employed in our study encompass a total of 415,523 galaxies, with an r-band absolute magnitude threshold set at -17.7. The sky distribution of these galaxies is illustrated by the black dots in Fig.1. Furthermore, the redshifts of these galaxies, measured in the CMB frame, span the range from $z=0.01$ to $z=0.1$, as depicted by the black bars in Fig.2. The galaxy mean number density $\bar{n}({\bf r})$ is shown in the top-left panel of Fig.3. The calculation of $\bar{n}({\bf r})$ is detailed in Bautista et al. (2025). In this paper, the density field is derived from the BGS catalogue.
@@ 2.1.2 The DESI-PV data
Due to the peculiar motions of galaxies, the apparent distance of a galaxy, denoted as $d_{z}$, deviates from its true comoving distance, $d_{h}$. This discrepancy is measureable and can be quantified through the logarithmic distance ratio, defined as $\eta\equiv\log_{10}\frac{d_{z}}{d_{h}}$. It can be converted to the line-of-sight peculiar velocity using the velocity estimator Eq.57. Within the DESI-PV survey, two distinct populations of galaxies are observed: spiral galaxies, also known as late-type galaxies, and elliptical galaxies, referred to as early-type galaxies. The $\eta$ for late-type galaxies is derived from the Tully-Fisher (TF, Tully & Fisher 1977) relation, whereas for early-type galaxies, it is inferred from the Fundamental Plane (FP, Djorgovski & Davis 1987, Dressler et al. 1987) relation. In DESI-PV, the Tully-Fisher sample (Douglass et al., 2025) comprises a total of 6,806 late-type galaxies. The sky distribution of these galaxies is illustrated by the blue dots in Fig.1. Furthermore, the redshifts of these galaxies span the range from $z=0.015$ to $z=0.1$, as depicted by the blue bars in Fig.2. In contrast, the Fundamental Plane sample (Ross et al., 2025) consists of 73,822 early-type galaxies. Their sky distribution is marked by gray dots in Fig.1, and their redshifts span the interval from $z=0.01$ to $z=0.1$, as shown in the gray bars of Fig.2.
@@ 2.2 Mocks
A considerable number of mocks is indispensable for validating the algorithm and accurately estimating the uncertainties in the power spectrum and associated parameters. Bautista et al. (2025) offers a comprehensive set of 675 BGS mocks, meticulously designed to mirror the characteristics of the BGS data. In parallel, 675 FP mocks have been constructed to emulate the FP data, alongside 675 TF mocks tailored to align with the TF data. Each set of BGS, FP, and TF mocks shares the same observer, resulting in a total of 675 mock sets. These mocks are derived from the AbacusSummit simulation, and the fiducial value of $f\sigma_{8}$ for these mocks is 0.466 at an effective redshift of $z=0.2$.
@@ 2.3 The random catalogs
The random catalogues for both the data and mocks are also supplied by Bautista et al. (2025). In order to precisely extract cosmological parameters from the power spectrum, it is imperative to appropriately account for the survey geometry of both BGS and DESI-PV. Employing fast Fourier transforms to measure the power spectrum in the context of a non-periodic and incomplete survey geometry leads to a convolution of the true power spectrum with the survey’s window function. To address this, the random catalogue will be employed to estimate the window function, thereby enabling the application of an equivalent convolution to our theoretical model during the data fitting process. The random catalogues for FP and TF are generated independently, with the objective of more faithfully replicating the distinct observational features of the TF and FP surveys, respectively. The random catalogue for DESI-PV is constructed by combining the random catalogues of TF and FP.
@@ 3.1 The field functions
We favor the momentum power spectrum over the velocity power spectrum because, in principle, the velocity field is a continuous entity defined everywhere in space—even in regions devoid of galaxies. However, our measurements are limited to locations where galaxies (and thus mass) are present, effectively yielding a mass-weighted, or momentum-based power spectrum. In regions devoid of galaxies, we record zero velocity not because the velocity is truly zero, but due to the absence of tracers. Additionally, the momentum power spectrum retains the same fundamental information as the velocity power spectrum, and with additional non-linear contributions introduced by the $(1+\delta)$ term that be incorporated into the theoretical modeling. There exist methods to probe the underlying continuous velocity field, however, they come with notable drawbacks. For example, the reconstruction of velocity field, it inherently dependent on cosmological assumptions. Alternatively, one may employ a Voronoi tessellation to discretize the data, thereby avoiding artificial zero-velocity measurements caused by sparse galaxy distribution. However, this method introduces complications in the Fourier space due to the irregularity of the resulting grid.
@@
which ensures that the amplitude of the measured power aligns with it in a universe unaffected by survey selection effects. The computation of the mean galaxy number density $\bar{n}_{\delta}({\bf r})$ at position $\bf r$} is detailed in Bautista et al. (2025) or see Fig.3. As indicated in Eq.6, to estimate the density field, it is necessary to subtract the random catalog, this is because we measure the density contrast field $\delta$, which can only be accomplished if we possess knowledge of, and are able to subtract, the mean number of galaxies within each grid cell which estimated from randoms of that cell.
@@
which again adjusts the amplitude of the measured power to conform with it in an idealized universe without survey selection biases. The computation of the mean galaxy number density $\bar{n}_{p}({\bf r})$ at position $\bf r$} is also detailed in Bautista et al. (2025) or see Fig.3. Notably, no random catalog is subtracted when estimating the momentum field. This is because that the mean velocity at any spatial point is inherently zero; consequently, subtracting a random catalogue during the construction of the momentum field would be equivalent to subtracting zero.
@@ 3.2 The estimators of the power spectrum multipoles
As outlined in Appendix H.1 (or see Paper III), the estimator for the cross-power spectrum multipoles is defined through  [^3] There is a small typos in Equation 19 of Paper III, while the code used in Paper III is correct.
@@
where $\ell=0,1,2,3,4,$.... Consequently, based on Eq.13 or Eq.14, the corresponding even-multipoles of the auto-power spectrum is expressed in (see Appendix H.3 for more discussion)
@@
where $d\Omega_{k}$ denotes the solid angle. See Appendix D for the expressions of $T_{0}$, $T_{2}$ and $T_{4}$. It is important to note that, unlike in Bianchi et al. (2015), the normalization factors $A_{\delta}$ and $A_{p}$ are inherently incorporated within the field functions in our formulation, and our definition of Fourier transformation diverges from them too. For multipoles with $\ell>0$, the shot noise becomes negligible and can be safely ignored (Blake et al., 2010; Howlett, 2019). If ignoring the wide-angle or general relativistic effects, only the even-multipoles of the auto-power spectrum are non-zero (Bianchi et al., 2015; Beutler & Di Dio, 2020; Castorina & White, 2020).
@@
The PYTHON code for our power spectrum estimators can be download from the website links provided in Appendix I or see https://github.com/FeiQin-cosmologist/Galaxy_Power_Spectrum/blob/main/CosmPSPy/Code/PSestFun.py.